# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 5. Recognition and uniqueness results for groups and modules
If $R\leq C_{V}(S)$, then $R\leq[W,S]=Z_{i-1}(S)$. In a similar manner to before, we apply Lemma 5.2 and Lemma 5.3 to deduce that $|W/R|=q^{b}$ for some $b\in\mathbb{N}$ and we conclude that either $R=Z(S)$ or $R=1$, and so we lie in the third case:
$Z_{0}\leq R=Z_{k}\leq Z_{i-2}<Z_{i}=R<Z_{n}=[V,S].$
@@
We have that $S/Z_{k}(S)\cong S_{n-k}(q)$ and by induction, using that $V/R$ is reducible, we deduce that $k=0$ and $W$ is irreducible. Applying Lemma 5.4 we deduce that $W\cong V_{i-1}(q)$. As $W<[V,S]$, $V/W\cong V_{n-i}(q)$ and by Proposition 5.7, using that $|V|=q^{n+1}$, we have a contradiction.
@@
We have shown that $i=n+1$ so that $[V,S]<W$ and $H$ centralises $V/W$. Then $[W,S]=[V,S]<W$ by Lemma 2.10. Let $k\in N_{H}(S)$ be of order $q-1$, set $K=\langle k\rangle$ and note that $\mathrm{Aut}_{K}(S)$ has order $q-1$, or $\frac{q-1}{2}$.
@@
Observe that $k$ induces an automorphism on $S$ which induces an action on $S/[V,S,S]$. But by Lemma 2.6, $S/[V,S,S]\cong S_{1}(q)$ is isomorphic to a Sylow $p$-subgroup of $\mathrm{PSL}_{3}(q)$. Since $q>p$, $k$ acts non-trivially on $S/V$ so acts non-trivially on $S/[V,S,S]$. By Proposition 2.5, $S/[V,S,S]$ is a $\mathbb{K}$-space for $\mathrm{Aut}(S/[V,S,S])$ and we conclude that $[k,V]\leq[V,S]$.
@@
We recognise that $\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$ contains a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S_{n}(q))$ and so for $K:=\langle k\rangle$, $\mathrm{Aut}_{K}(S)$ is $\mathrm{Aut}(S)$-conjugate to an element of order at least $\frac{q-1}{2}$ which acts like an element of $\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$. Since $q>p$, Lemma 2.7 implies that $K$ acts irreducibly on $C_{V}(S)=Z_{1}$ as an $\mathbb{F}_{p}K$-module.
@@
Let $1\neq Q\leq W<V$ with $Q$ chosen minimally such that $Q\trianglelefteq H$. In particular, $Q$ is an irreducible $H/V$-module. Since $K$ acts irreducibly on $C_{V}(S)$, we see that $C_{V}(S)=\langle(C_{V}(S)\cap Q)^{K}\rangle\leq Q$. As before, applying Lemma 5.2 and Lemma 5.3 we conclude that $Q$ is writable as a $\mathbb{K}$-module and so $Q=Z_{i}(S)$ for some $1\leqslant i\leqslant n$ by Lemma 5.4. If $i<n$, then $S/Q\cong S_{n-i}(q)$ and by induction we have that $V/Q$ is irreducible which is impossible as $Q<W$. Hence $Q=[V,S]\cong V_{n-1}(q)|_{\mathbb{F}_{p}}$ and $V/Q$ is centralised by $H$. An application of Lemma 5.8(2) then yields the unforgivable $[V,S]=[V,S,S]$. This completes the proof. ∎
@@
We observe that $Z(S)=C_{V}(S)$ has order $q^{2}$ and $V/[V,S]$ has order $q$ by Lemma 2.12. Let $X\leq V$ be such that $X$ corresponds to the $(p-1)$-dimensional $\mathbb{K}\operatorname{SL}_{2}(q)$-submodule in the construction of $S_{\Lambda}(q)$. In particular, $S/X\cong S_{1}(q)$, $[X,S]=[V,S,S]$, $C_{X}(S)=C_{[V,S,S]}(S)$ and $[V,S]=XC_{V}(S)$.
As in Proposition 5.14, we have that $H/C_{H}(V)\cong\operatorname{SL}_{2}(q)$ or $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$ and we let $k\in N_{H}(S)$ be of order $q-1$ so that $\mathrm{Aut}_{\langle k\rangle}(S)$ has order $q-1$ or $\frac{q-1}{2}$ respectively. Write $K:=\langle k\rangle$. Then $\mathrm{Aut}_{K}(S)$ is an abelian subgroup of $\mathrm{Aut}(S)$ of order at least $\frac{q-1}{2}$. Since $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ contains a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S_{\Lambda}(q))$, $\mathrm{Aut}_{K}(S)$ is conjugate in $\mathrm{Aut}(S)$ to a subgroup of the unique abelian subgroup of order $(q-1)^{2}$ in a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ by Lemma 2.15.
Let $tV\in H/V\cong\operatorname{SL}_{2}(q)$ with $t$ an involution. In particular, we may arrange that $t$ is the unique involution in $K$. Then $C_{V}(t)\geq C_{V}(H)\cap C_{X}(S)\neq 1$ and $V=[V,t]\times C_{V}(t)$ is an $H$-invariant decomposition of $V$. If $[V,t]\not\leq[V,S]$, then $[V,t]\geq[V,S]$ and if $C_{V}(t)\not\leq[V,S]$, then $C_{V}(t)\geq[V,S]$. Since $C_{V}(t)\cap[V,t]=1$, we deduce that either $V=[V,t]$ or $V=C_{V}(t)$.
@@
Since $K$ centralises $C_{V}(H)\cap C_{X}(S)$, and a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{X}(S)$, we deduce that $K$ centralises $C_{X}(S)$. It follows that $C_{V}(K)=C_{X}(S)$ by Lemma 2.16. Then, as $C_{V}(t)\neq 1$, we also know that $C_{V}(t)=V$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$.
@@
Let $Y\trianglelefteq H$ with $Y<V$, and choose $Y$ maximally with respect to these conditions. Then $V/Y$ is an irreducible module for $H/V$. If $Y\not<[V,S]$ then $[V,S]\leq Y$ so that $H$ centralises $V/Y$. Indeed, as $Y<V$ and $K$ centralises $V/Y$, we must have that $C_{S}(K)\neq C_{X}(S)$, a contradiction. Hence, $Y<[V,S]$ and so $[V/Y,S]$ has index $q$ in $V/Y$. Applying Lemma 5.2, we see that $|V/Y|=q^{a}$, for some $a\in\mathbb{N}$ and $C_{V/Y}(S)$ has order $q$. Furthermore, both $C_{V/Y}(S)$ and $V/[V,S]$ are irreducible as $\mathbb{F}_{p}K$-modules. Since $C_{V/Y}(S)$ is irreducible as a $\mathbb{F}_{p}K$-module, we must have that either $C_{V}(S)\leq Y$ or $C_{V/Y}(S)=C_{V}(S)Y/Y$.
@@
Assume first that $C_{V}(S)\leq Y$. Then, as $S/C_{V}(S)\cong S_{p-2}(q)$ by Lemmas 2.8 and 2.14, we have that $S/Y\cong S_{a-1}(q)$. Since $V/Y$ is irreducible, Proposition 5.14 implies that $V/Y\cong V_{a-1}(q)|_{\mathbb{F}_{p}}$ as an $H/C_{V}(V/Y)$-module. Since $t$ centralises $V$, $a-1$ is even. But then, $K$ centralises $[V/Y,S;\frac{a-1}{2}]/[V/Y,S;\frac{a+1}{2}]$, contrary to $C_{V}(K)=C_{X}(S)$. Hence $C_{V}(S)\not\leq Y$.
@@
Now, again applying Lemmas 2.8 and 2.14, we have that $S/C_{V}(S)Y\cong S_{a-2}(q)$ and it follows that $[V/Y,S;a]=1$ but $[V/Y,S;a-1]\neq 1$. Appealing to Lemma 5.4, we ascertain that $V/Y\cong V_{a-1}(q)|_{\mathbb{F}_{p}}$ and as $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$, we again see that $a-1$ is even. As above, this implies that $C_{V}(K)>C_{X}(S)$, and we have a contradiction by Lemma 2.16.
Case 2: $C_{V}(H)\cap C_{X}(S)=1$.
@@
Suppose $Y$ satisfies $Y\trianglelefteq H$, $Y<V$ and $[H,Y]\neq 1$ and $Y$ is chosen minimally with respect to these conditions. In particular, $Y/C_{Y}(H)$ is irreducible. Note that $C_{Y}(H)\cap C_{X}(S)=1$. If $Y\not\leq C_{V}(S)[V,S,S]$ then $[Y,S]=[V,S,S]$ so that $C_{X}(S)\leq Y$. On the other hand, if $Y\leq C_{V}(S)[V,S,S]$ then $[Y,S]\leq[V,S,S]$. Indeed, since $[Y,S]\neq 1$, we must have that $C_{X}(S)\leq[Y,S]$. In either case, applying Lemma 5.2 we conclude that $K$ acts irreducibly on $C_{X}(S)$ and $|Y/C_{Y}(H)|=q^{a}$ for some $a\in\mathbb{N}$.
@@
Assume first that $C_{Y}(H)\neq 1$ so that $V=C_{V}(t)$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{V}(S)[V,S,S]/[V,S,S]$, in combination with Lemma 2.16 we get that $C_{V}(S)=C_{V}(K)\times C_{X}(S)$ and $|C_{V}(K)|=q$.
@@
If $Y\not\leq C_{V}(S)[V,S,S]$, then as $K$ acts irreducibly on $Y/C_{Y}(H)[Y,S]$ and $|Y/C_{Y}(H)|=q^{a}$, we deduce that $|V/Y|<q^{2}$ so that $H$ centralises $V/Y$. Since $|C_{V}(K)|=q$, the only possibility is that $V=C_{V}(S)Y$ and as $C_{V}(S)\leq[V,S]$, we conclude that $V=Y$. Then $C_{V}(K)=C_{V}(H)$ and applying Lemma 5.4, we have that $V/C_{V}(H)\cong V_{p-1}(q)|_{\mathbb{F}_{p}}$. But $K$ centralises $[V/C_{V}(H),S;\frac{p-1}{2}]/[V/C_{V}(H),S;\frac{p+1}{2}]$, contrary to $|C_{V}(K)|=q$.
@@
If $Y\leq C_{V}(S)[V,S,S]$ then since $C_{V}(H)\cap C_{X}(S)=1$ and $|C_{Y/C_{Y}(H)}(S)|=q$, we deduce that $Y=(Y\cap[V,S,S])C_{Y}(H)$. Since $|Y/[Y,S]C_{Y}(H)|=q$, we deduce that $Y\cap[V,S,S]=[V,S;i]$ for some $i\in\mathbb{N}$. Then Lemma 5.4 reveals that $Y/C_{Y}(H)\cong V_{p-i-1}(q)|_{\mathbb{F}_{p}}$ where $i$ is even. But then $K$ centralises $[Y/C_{Y}(H),S;\frac{p-i-1}{2}]/[Y/C_{Y}(H),S;\frac{p-i+1}{2}]$, against $|C_{V}(K)|=q$.
@@
Hence $C_{Y}(H)=1$ and $Y$ is irreducible as an $\mathbb{F}_{p}H$-module. In particular, $|Y|\geqslant q^{2}$ and $C_{Y}(S)=C_{X}(S)$. Observe that if $Y\not\leq[V,S]$ then $[Y,S]=[V,S]$ so that $|C_{Y}(S)|=q^{2}$, a contradiction. Furthermore, if $Y\not\leq C_{V}(S)[V,S,S]$, then $[Y,S]=[V,S,S]$ and as $|Y/[Y,S]|=q$ and $C_{Y}(S)=C_{X}(S)$, we see by Lemma 5.4 that $Y\cong V_{p-2}(q)|_{\mathbb{F}_{p}}$ and $|V/Y|=q^{2}$. If $[K,V]\leq Y$ then $[K,C_{V}(S)]\leq C_{X}(S)$, $|C_{V}(K)|>q$ and Lemma 2.16 gives a contradiction. Hence, $V/Y\cong V_{1}(q)|_{\mathbb{F}_{p}}$ and Proposition 5.7 implies that $V\cong\Lambda(q)|_{\mathbb{F}_{p}}$, as desired.
@@
Thus, for the remainder of the proof we assume, aiming for a contradiction, that $Y\leq C_{V}(S)[V,S,S]$. Let $i\in\mathbb{N}$ be such that $Y\leq C_{V}(S)[V,S;i]$ but $Y\not\leq C_{V}(S)[V,S;i+1]$. Then $[Y,S]=[V,S;i+1]$ and since $|Y|=q^{a}$ for some $a\in\mathbb{N}$, we deduce that $|Y/[Y,S]|=q$, $a=p-i$ and $Y\cong V_{p-i-1}(q)|_{\mathbb{F}_{p}}$. Set $\overline{V}=V/[V,S;i+1]$. Then $K$ acts irreducibly on $\overline{Y}$ and normalises $\overline{[V,S;i]}$ and $\overline{C_{V}(S)}$. Moreover, $\overline{Y}\cap\overline{C_{V}(S)}=1$.
@@
Hence, $Y=[V,S;i]$. Let $Q\trianglelefteq H$ such that $Y<Q\leq V$ and $Q$ is minimal subject to these conditions. Then $H$ acts irreducibly on $Q/Y$. If $[Q,S]\leq Y$ then $[Q,H]\leq Y$ and we deduce from Lemma 5.8 that $[Q,S]=[Y,S]$ and $Q\leq C_{V}(S)Y$. Then $[K,W]\leq Y$, $V=C_{V}(t)$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$ so that $a$ is odd. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{V}(S)[V,S,S]/[V,S,S]$, in combination with Lemma 2.16 we get that $C_{V}(S)=C_{V}(K)\times C_{X}(S)$ and $|C_{V}(K)|=q$. But $K$ also centralises $[Y,S;\frac{p-i-1}{2}]/[Y,S;\frac{p-i+1}{2}]$, against $|C_{V}(K)|=q$. Hence, $[Q,S]\not\leq Y$ from which we deduce that $[V,S;i-1]\leq Q$ and $[V,S;i-1]/Y=C_{Q/Y}(S)$ by Lemmas 5.2 and 5.3.
From Lemma 5.2 we see that $|Q/Y|=q^{a}$ for some $a\in\mathbb{N}$. If $Q\not\leq[V,S]$ then $C_{V}(S)[V,S;i-1]/Y=C_{Q/Y}(S)$ has order $q^{2}$, a contradiction by Lemma 5.3. Then $Q\leq[V,S]$. Since $K$ acts irreducibly on $Q/[Q,S]$, we deduce that either $Q\leq[V,S,S]$ or $[Q,S]=Q\cap[V,S,S]$. In either instance, $Q\cap[V,S,S]=[V,S;j]$ for some $j\in\mathbb{N}$ and Lemma 5.4 implies that $Q/Y\cong V_{j-i-1}(q)|_{\mathbb{F}_{p}}$, against Proposition 5.7. This completes the proof. ∎

### v2

@@ 5. Recognition and uniqueness results for groups and modules
If $W_{0}\leq C_{V}(S)$, then $W_{0}\leq[W,S]=Z_{i-1}$. In a similar manner to before, we apply Lemma 5.2 and Lemma 5.3 to deduce that $|W/W_{0}|=q^{b}$ for some $b\in\mathbb{N}$ and we conclude that either $W_{0}=Z(S)$ or $W_{0}=1$, and so $Z_{k}$ for some $k$. Thus, we lie in the third case:
$1\leq W_{0}=Z_{k}\leq Z_{i-2}<Z_{i}=W<Z_{n}=[V,S].$
@@
We have that $S/Z_{k}\cong S_{n-k}(q)$ and by induction, using that $V/W_{0}$ is reducible, we deduce that $k=0$ and $W$ is irreducible. Applying Lemma 5.4 we deduce that $W\cong V_{i-1}(q)$. As $W<[V,S]$, $V/W\cong V_{n-i}(q)$ and by Proposition 5.7, using that $|V|=q^{n+1}<q^{p+1}$, we have a contradiction. This proves the claim.
@@
We have shown that $i=n+1$ so that $[V,S]<W$ and $H$ centralises $V/W$. Then $[W,S]=[V,S]<W$ by Lemma 2.14. Let $k\in N_{H}(S)$ be of order q1q-1, set $K=\langle k\rangle$ and note that $\mathrm{Aut}_{K}(S)$ has order q1q-1, or $\frac{q-1}{2}$.
@@
Observe that $k$ induces an automorphism on $S$ which induces an action on $S/[V,S,S]$. But by Lemma 2.10, $S/[V,S,S]\cong S_{1}(q)$ is isomorphic to a Sylow $p$-subgroup of $\mathrm{PSL}_{3}(q)$. Since $q>p$, $k$ acts non-trivially on $S/V$ so acts non-trivially on $S/[V,S,S]$. By Proposition 2.9, $S/[V,S,S]$ is a $\mathbb{K}$-space for $\mathrm{Aut}(S/[V,S,S])$ and we conclude that $k,V]\leq[V,S]$.
@@
Recall the definition of $P_{n}^{*}(q)$ from Notation 2.3. We recognise that $\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$ contains a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S_{n}(q))$ and so for $K:=\langle k\rangle$, $\mathrm{Aut}_{K}(S)$ is $\mathrm{Aut}(S)$-conjugate to an element of order at least $\frac{q-1}{2}$ which acts like an element of $\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$. Since $q>p$, Lemma 2.11 implies that $K$ acts irreducibly on $C_{V}(S)=Z_{1}$ as an $\mathbb{F}_{p}K$-module.
@@
Let $1\neq V_{0}\leq W<V$ with $V_{0}$ chosen minimally such that $V_{0}\trianglelefteq H$. In particular, $V_{0}$ is an irreducible $H/V$-module. Since $K$ acts irreducibly on $C_{V}(S)$, we see that $C_{V}(S)=\langle(C_{V}(S)\cap V_{0})^{K}\rangle\leq V_{0}$. As before, applying Lemma 5.2 and Lemma 5.3 we conclude that $V_{0}$ is writable as a $\mathbb{K}$-module and so $V_{0}=Z_{i}(S)$ for some $1\leqslant i\leqslant n$ by Lemma 5.4. If $i<n$, then $S/V_{0}\cong S_{n-i}(q)$ and by induction we have that $V/V_{0}$ is irreducible which is impossible as $V_{0}<W$. Hence $V_{0}=[V,S]\cong V_{n-1}(q)|_{\mathbb{F}_{p}}$ and $V/V_{0}$ is centralised by $H$. An application of Lemma 5.8(2) then yields the unforgivable $[V,S]=[V,S,S]$. This completes the proof. ∎
@@
We observe that $Z(S)=C_{V}(S)$ has order $q^{2}$ and $V/[V,S]$ has order $q$ by Lemma 2.16. Let $W\leq V$ be such that $W$ corresponds to the (p1)(p-1)-dimensional $\mathbb{K}\operatorname{SL}_{2}(q)$-submodule in the construction of $S_{\Lambda}(q)$. In particular, $S/W\cong S_{1}(q)$, $[W,S]=[V,S,S]$, $C_{W}(S)=C_{[V,S,S]}(S)$ and $V,S]=WC_{V}(S)$.
As in Proposition 5.14, we have that $H/C_{H}(V)\cong\operatorname{SL}_{2}(q)$ or $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$ and we let $k\in N_{H}(S)$ be of order q1q-1 so that $\mathrm{Aut}_{\langle k\rangle}(S)$ has order q1q-1 or $\frac{q-1}{2}$ respectively. Write $K:=\langle k\rangle$. Then $\mathrm{Aut}_{K}(S)$ is an abelian subgroup of $\mathrm{Aut}(S)$ of order at least $\frac{q-1}{2}$. Recall the definition of $P_{\Lambda}^{*}(q)$ from Notation 2.5. Since $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ contains a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S_{\Lambda}(q))$, $\mathrm{Aut}_{K}(S)$ is conjugate in $\mathrm{Aut}(S)$ to a subgroup of the unique abelian subgroup of order $(q-1)^{2}$ in a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ by Lemma 2.19.
Let $tV\in H/V\cong\operatorname{SL}_{2}(q)$ with $t$ an involution. In particular, we may arrange that $t$ is the unique involution in $K$. Then $C_{V}(t)\geq C_{V}(H)\cap C_{W}(S)\neq 1$ and $V=[V,t]\times C_{V}(t)$ is an $H$-invariant decomposition of $V$. If $V,t]\not\leq[V,S]$, then $V,t]\geq[V,S]$ and if $C_{V}(t)\not\leq[V,S]$, then $C_{V}(t)\geq[V,S]$. Since $C_{V}(t)\cap[V,t]=1$, we deduce that either $V=[V,t]$ or $V=C_{V}(t)$.
@@
Since $K$ centralises $C_{V}(H)\cap C_{W}(S)$, and a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{W}(S)$, we deduce that $K$ centralises $C_{W}(S)$. It follows that $C_{V}(K)=C_{W}(S)$ by Lemma 2.20. Then, as $C_{V}(t)\neq 1$, we also know that $C_{V}(t)=V$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$.
@@
Let $V_{M}\trianglelefteq H$ with $V_{M}<V$, and choose $V_{M}$ maximally with respect to these conditions. Then $V/V_{M}$ is an irreducible module for $H/V$. If $V_{M}\not<[V,S]$ then $V,S]\leq V_{M}$ so that $H$ centralises $V/V_{M}$. Indeed, as $V_{M}<V$ and $K$ centralises $V/V_{M}$, we must have that $C_{V}(K)\neq C_{W}(S)$, a contradiction. Hence, $V_{M}<[V,S]$ and so $V/V_{M},S]$ has index $q$ in $V/V_{M}$. Applying Lemma 5.2, we see that $|V/V_{M}|=q^{a}$, for some $a\in\mathbb{N}$ and $C_{V/V_{M}}(S)$ has order $q$. Furthermore, both $C_{V/V_{M}}(S)$ and $V/[V,S]$ are irreducible as $\mathbb{F}_{p}K$-modules. Since $C_{V/V_{M}}(S)$ is irreducible as a $\mathbb{F}_{p}K$-module, we must have that either $C_{V}(S)\leq V_{M}$ or $C_{V/V_{M}}(S)=C_{V}(S)V_{M}/V_{M}$.
@@
Assume first that $C_{V}(S)\leq V_{M}$. Then, as $S/C_{V}(S)\cong S_{p-2}(q)$ by Lemmas 2.12 and 2.18, we have that $S/V_{M}\cong S_{a-1}(q)$. Since $V/V_{M}$ is irreducible, Proposition 5.14 implies that $V/V_{M}\cong V_{a-1}(q)|_{\mathbb{F}_{p}}$ as an $H/C_{H}(V/V_{M})$-module. Since $t$ centralises $V$, a1a-1 is even. But then, $K$ centralises $V/V_{M},S;\frac{a-1}{2}]/[V/V_{M},S;\frac{a+1}{2}]$, contrary to $C_{V}(K)=C_{W}(S)$. Hence $C_{V}(S)\not\leq V_{M}$.
@@
Now, again applying Lemmas 2.12 and 2.18, we have that $S/C_{V}(S)V_{M}\cong S_{a-2}(q)$ and it follows that $V/V_{M},S;a]=1$ but $V/V_{M},S;a-1]\neq 1$. Appealing to Lemma 5.4, we ascertain that $V/V_{M}\cong V_{a-1}(q)|_{\mathbb{F}_{p}}$ and as $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$, we again see that a1a-1 is even. As above, this implies that $C_{V}(K)>C_{W}(S)$, and we have a contradiction by Lemma 2.20.
Case 2: $C_{V}(H)\cap C_{W}(S)=1$.
@@
Suppose $Y$ satisfies $Y\trianglelefteq H$, $Y<V$ and $H,Y]\neq 1$ and $Y$ is chosen minimally with respect to these conditions. In particular, $Y/C_{Y}(H)$ is irreducible. Note that $C_{Y}(H)\cap C_{W}(S)=1$. If $Y\not\leq C_{V}(S)[V,S,S]$ then $[Y,S]=[V,S,S]$ so that $C_{W}(S)\leq Y$. On the other hand, if $Y\leq C_{V}(S)[V,S,S]$ then $Y,S]\leq[V,S,S]$. Indeed, since $Y,S]\neq 1$, we must have that $C_{W}(S)\leq[Y,S]$. In either case, applying Lemma 5.2 we conclude that $K$ acts irreducibly on $C_{W}(S)$ and $|Y/C_{Y}(H)|=q^{a}$ for some $a\in\mathbb{N}$.
@@
Assume first that $C_{Y}(H)\neq 1$ so that $V=C_{V}(t)$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{V}(S)[V,S,S]/[V,S,S]$, in combination with Lemma 2.20 we get that $C_{V}(S)=C_{V}(K)\times C_{W}(S)$ and $|C_{V}(K)|=q$.
@@
If $Y\not\leq C_{V}(S)[V,S,S]$, then as $K$ acts irreducibly on $Y/C_{Y}(H)[Y,S]$ and $|Y/C_{Y}(H)|=q^{a}$, we deduce that $|V/Y|<q^{2}$ so that $H$ centralises $V/Y$. Since $|C_{V}(K)|=q$, the only possibility is that $V=C_{V}(S)Y$ and as $C_{V}(S)\leq[V,S]$, we conclude that $V=Y$. Then $C_{V}(K)=C_{V}(H)$ and applying Lemma 5.4, we have that $V/C_{V}(H)\cong V_{p-1}(q)|_{\mathbb{F}_{p}}$. But $K$ centralises $V/C_{V}(H),S;\frac{p-1}{2}]/[V/C_{V}(H),S;\frac{p+1}{2}]$, contrary to $|C_{V}(K)|=q$.
@@
If $Y\leq C_{V}(S)[V,S,S]$ then since $C_{V}(H)\cap C_{W}(S)=1$ and $|C_{Y/C_{Y}(H)}(S)|=q$, we deduce that $Y=(Y\cap[V,S,S])C_{Y}(H)$. Since $|Y/[Y,S]C_{Y}(H)|=q$, we deduce that $Y\cap[V,S,S]=[V,S;i]$ for some $i\in\mathbb{N}$. Then Lemma 5.4 reveals that $Y/C_{Y}(H)\cong V_{p-i-1}(q)|_{\mathbb{F}_{p}}$ where $i$ is even. But then $K$ centralises $Y/C_{Y}(H),S;\frac{p-i-1}{2}]/[Y/C_{Y}(H),S;\frac{p-i+1}{2}]$, against $|C_{V}(K)|=q$.
@@
Hence $C_{Y}(H)=1$ and $Y$ is irreducible as an $\mathbb{F}_{p}H$-module. In particular, $|Y|\geqslant q^{2}$ and $C_{Y}(S)=C_{W}(S)$. Observe that if $Y\not\leq[V,S]$ then $[Y,S]=[V,S]$ so that $|C_{Y}(S)|=q^{2}$, a contradiction. Furthermore, if $Y\not\leq C_{V}(S)[V,S,S]$, then $[Y,S]=[V,S,S]$ and as $|Y/[Y,S]|=q$ and $C_{Y}(S)=C_{W}(S)$, we see by Lemma 5.4 that $Y\cong V_{p-2}(q)|_{\mathbb{F}_{p}}$ and $|V/Y|=q^{2}$. If $K,V]\leq Y$ then $K,C_{V}(S)]\leq C_{W}(S)$, $|C_{V}(K)|>q$ and Lemma 2.20 gives a contradiction. Hence, $V/Y\cong V_{1}(q)|_{\mathbb{F}_{p}}$ and Proposition 5.7 implies that $V\cong\Lambda(q)|_{\mathbb{F}_{p}}$, as desired.
@@
Thus, for the remainder of the proof we assume, aiming for a contradiction, that $Y\leq C_{V}(S)[V,S,S]$. Let $i\in\mathbb{N}$ be such that $Y\leq C_{V}(S)[V,S;i]$ but $Y\not\leq C_{V}(S)[V,S;i+1]$. Then $[Y,S]=[V,S;i+1]$ and since $|Y|=q^{a}$ for some $a\in\mathbb{N}$, we deduce that $|Y/[Y,S]|=q$, a=pia=p-i and $Y\cong V_{p-i-1}(q)|_{\mathbb{F}_{p}}$. Set $\overline{V}=V/[V,S;i+1]$. Then $K$ acts irreducibly on $\overline{Y}$ and normalises $\overline{[V,S;i]}$ and $\overline{C_{V}(S)}$. Moreover, $\overline{Y}\cap\overline{C_{V}(S)}=1$.
@@
Hence, $Y=[V,S;i]$. Let $V_{0}\trianglelefteq H$ such that $Y<V_{0}\leq V$ and $V_{0}$ is minimal subject to these conditions. Then $H$ acts irreducibly on $V_{0}/Y$. If $V_{0},S]\leq Y$ then $V_{0},H]\leq Y$ and we deduce from Lemma 5.8 that $V_{0},S]=[Y,S]$ and $V_{0}\leq C_{V}(S)Y$. Then $K,V_{0}]\leq Y$, $V=C_{V}(t)$ and $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$ so that $a$ is odd. Since a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}_{P_{\Lambda}^{*}(q)}(S_{\Lambda}(q))$ acts irreducibly on $C_{V}(S)[V,S,S]/[V,S,S]$, in combination with Lemma 2.20 we get that $C_{V}(S)=C_{V}(K)\times C_{W}(S)$ and $|C_{V}(K)|=q$. But $K$ also centralises $Y,S;\frac{p-i-1}{2}]/[Y,S;\frac{p-i+1}{2}]$, against $|C_{V}(K)|=q$. Hence, $V_{0},S]\not\leq Y$ from which we deduce that $V,S;i-1]\leq V_{0}$ and $V,S;i-1]/Y=C_{V_{0}/Y}(S)$ by Lemmas 5.2 and 5.3.
From Lemma 5.2 we see that $|V_{0}/Y|=q^{a}$ for some $a\in\mathbb{N}$. If $V_{0}\not\leq[V,S]$ then $C_{V}(S)[V,S;i-1]/Y=C_{V_{0}/Y}(S)$ has order $q^{2}$, a contradiction by Lemma 5.3. Then $V_{0}\leq[V,S]$. Since $K$ acts irreducibly on $V_{0}/[V_{0},S]$, we deduce that either $V_{0}\leq[V,S,S]$ or $V_{0},S]=V_{0}\cap[V,S,S]$. In either instance, $V_{0}\cap[V,S,S]=[V,S;j]$ for some $j\in\mathbb{N}$ and Lemma 5.4 implies that $V_{0}/Y\cong V_{j-i-1}(q)|_{\mathbb{F}_{p}}$, against Proposition 5.7. This completes the proof. ∎